Consider the following setup for a hypothetical economy:

People Employed = 1,573,825
Labor Force = 1,984,327
Total Population = 2,341,192
Natural Rate of Unemployment = 5.65%
Discouraged Workers = 238,649
------------------------------------------------------------

How many people are either structurally or frictionally unemployed?

Note: Enter the actual number of people.
1
Expert's answer
2011-10-20T12:58:25-0400
Total unemployment = (1984327-1573825)/2341192 = 0.1753 = 17.53%, %discouraged =
238649/2341192 = 10.19%,
%structurally and frictionally unemployed =
17.53%-10.19%-5.65% = 1.69%, structurally and frictionally unemployed = 1.69%
2341192 = 16154 people.